# Funny Valen-Tine: Solving visual abstract reasoning problems through defining the solution distribution

### v1

@@ Abstract
The importance of visual abstruct reasoning problems in the field of image processing is beyond words. Both Bongard-Logo problems and Raven’s Progressive Matrices (RPM) belong to the domain of visual abstract reasoning tasks, with Bongard-Logo categorized as image clustering reasoning and RPM involving image progression pattern reasoning. This paper introduces a novel baseline model, Valen, which falls under the umbrella of probabilistic highlighting models. Valen demonstrates remarkable performance in solving both RPM and Bongard-Logo problems, offering a versatile solution for these reasoning tasks. Our investigation extends beyond the application of Valen, delving into the underlying mechanisms of probability-highlighting solvers when tackling visual abstract reasoning problems. We revisit how these solvers handle instances in tasks like RPM and Bongard-Logo, realizing that they approximate the solution to each reasoning problem instance as a distribution delineated by primary and auxiliary samples. This led us to propose that the learning objective of probability-highlighting solvers is not the distribution of correct solutions but rather a distribution defined by both primary and auxiliary samples. To bridge the discrepancies, we introduced the Tine method, an adversarial learning-based approach designed to assist Valen in estimating a solution distribution closer to the correct one. However, Tine’s adversarial framework also introduces issues such as unstable training. Reflecting on Tine, we propose modeling the sample distribution of reasoning problems as a mixture of Gaussian distributions, leading to the development of the Funny method. Funny effectively models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution. Furthermore, we designed the SBR method to model the distribution of progressive patterns representation as a mixture of Gaussian distributions. Overall, the Funny, Tine, and SBR methods have significantly improved Valen’s performance, providing new ideas and methods for the study of visual abstract reasoning problems.
@@
Abstract reasoning, Raven’s Progressive Matrices, Bongard-Logo problem, Adversarial Learning, Gaussian Mixture Model.
@@ I-A RAVEN and PGM
RPM[14], or Raven’s Progressive Matrices, is a classic visual reasoning intelligence test typically presented as a 3x3 matrix. It requires test-takers to observe and analyze visual features such as shape, size, direction, position, and their relationships within the problem statement, identify patterns or rules, and select an answer from given option pool that completes the matrix. Due to its non-verbal nature and value in assessing general intelligence, RPM is widely used in psychology, education, neuroscience, and other fields, serving as a tool for evaluating intelligence levels and studying human cognitive processes, intellectual development, and brain function. With the advancement of AI technology, researchers are exploring computer algorithms to solve RPM problems, albeit facing challenges.
RAVEN[15] and PGM[16] are classic examples of RPM problems, with Figure 1 illustrating these two types of problems respectively.
@@ II-A RPM solvers and Bongard-Logo solvers
There are numerous excellent solvers for RPM (Raven’s Progressive Matrices) and Bongard-Logo problems, among which CoPINet[19], LEN+teacher[18], DCNet[20], NCD[22], SCL[23], SAVIR-T[24], PrAE[25], NVSA[27], ALANS[26], RS-TRAN[28] and CRAB[29] are notable RPM solvers, while SBPDL[30], PMoC[31] and Triple-CFN[32] are outstanding Bongard-Logo solvers. Since the method proposed in this paper is relatively independent and does not have a strong dependency on these existing works, the details about them will not be elaborated further.
@@ III Methodology
This paper argues that explicitly defining the distribution of solutions for the designed abstract reasoning problem solver is a straightforward and effective approach. To this end, we first designed a general solver, Valen, for both RPM and Bongard-Logo problems. Additionally, we developed three accompanying methods: Funny, Tine, and SBR. These three methods all aim to explicitly define the distribution of correct solutions for Valen through explicit or implicit approaches.
Here we clarify a definition: in this paper, samples that conform to the underlying patterns within instances of reasoning problems are referred to as “primary samples.” For example, the correct options within RPM problem instances and the samples in the primary group within Bongard-Logo problem instances are considered primary samples. Conversely, samples that do not conform to the underlying patterns, such as the incorrect options in RPM problem instances and the samples in the auxiliary group within Bongard-Logo problem instances, are referred to as “auxiliary samples.”
@@ IV Valen: a new baseline model for RPM and Bongard-Logo problems
In this section, we introduce Valen (Visual Abstraction Learning Network), our baseline model designed to address both RPM (Raven’s Progressive Matrices), an image progressive pattern reasoning problem, and Bongard-Logo, an image clustering reasoning problem.
@@ IV-A Valen targeting RPM problem
The RPM problem challenges participants to complete a 3×3 matrix of images by choosing answers from a pool of options. A correct answer is one that, when placed in the matrix, maintains a consistent and specific image progression pattern that can be deduced from the incomplete matrix, referred to as the problem statement. [39].
@@ IV-A1 Representation extraction of matrix and enumeration of incomplete matrix
Subsequently, these multi-viewpoint representations $\{r_{i}|i\in[1,9]\}_{v}|_{v=1}^{V}$ will be enumerated by Valen to generate all possible incomplete representations matrix. The enumeration results are illustrated in the Figure 5, with the last item of the enumeration is the matrix with a incomplete option position.
@@
We denote these enumerated multi-viewpoint incomplete representation matrices as $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$, where $n$ represents the index of the incomplete representation matrix, and $i$ still represents the number of the representation within the incomplete matrix. Please note the order of the enumeration results set in the figure.
@@ IV-A2 Extraction of progressive patterns from incomplete matrix
The figure illustrates the process of the Transformer-Encoder handling $\{r_{9i}|i\in[1,8]\}_{v}$, which represents the last item of the enumeration results of incomplete representation matrices from a specific viewpoint $v$. The process of handling other incomplete representation matrices $\{r_{ni}|n\in[1,8]i\in[1,9],i\neq n\}_{v}$ is similar to the one shown in the figure, with no essential differences.
@@ V Tine, a Method for Advancing the Solution Distribution Boundary
Therefore, we can interpret these solvers as a conditional probability distributions obeyed by abstract reasoning problems, or understand them as providing distributions of solutions to abstract reasoning problems. If we use RPM solvers as examples, then these solvers can be abstracted as a conditional probability distribution, denoted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$. (Accordingly, when Valen is applied to Bongard-Logo problems, it can be abstracted as a conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{6},\theta)$). In this context, the two conditions of the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ are the statement of the reasoning problem $\{x_{i}\}_{i=1}^{8}$ and the optimizable parameters $\theta$ of the solver. Here, the $x_{\alpha}$ refers to the option that is pending verification. This conditional distribution implies that, when given a problem statement $\{x_{i}\}_{i=1}^{8}$ and an option $x_{\alpha}$, it is possible to calculate the probability of $x_{\alpha}$ being a solution to $\{x_{i}\}_{i=1}^{8}$.
@@
The optimization of probability-highlighting solver can be regarded as optimizing the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$. A wide range of loss functions can be used for optimization, but they all share the same objective: to assign high probabilities to the primary samples in the probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ and low probabilities to the auxiliary options. Such an optimization objective makes us realize that the optimization process of the probability-highlighting solver is not aimed at fitting the distribution of correct solutions to visual abstract reasoning problems, but rather is dedicated to fitting a distribution jointly delimited by primary and auxiliary samples.
@@
To simplify the explanation of our findings, we assume that the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ follows a Gaussian distribution. Consequently, we can explicitly identify the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta})$. Then, the negative logarithm of the probability density function of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta})$ can be expressed as follows:
$-\log(P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta}))=\frac{{(x_{a}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(|\,\sigma_{\theta}|\,)$ (1)
@@
Evidently, we have omitted the constant term $-1/2\log(2\pi)$. In this formula, $\mu_{\theta}$ and $\sigma_{\theta}$ respectively represent the mean and standard deviation of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$. Since we have defined the form of distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a Gaussian distribution, we consequently abstract the optimizable parameters $\theta$ into optimizable mean $\mu_{\theta}$ and standard deviation $\sigma_{\theta}$.
@@
Optimizing the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\mu_{\theta},\sigma_{\theta})$ using the loss function is equivalent to optimizing the values of $\mu_{\theta}$ and $\sigma_{\theta}$. Our optimization objective is that when primary samples (the correct options in an RPM problem, or the samples within the main group in a Bongard-Logo problem) are plugged into formula (1), we expect to obtain a smaller value. Conversely, when auxiliary samples (the incorrect options in an RPM problem, or the samples within the auxiliary group in a Bongard-Logo problem) are plugged into formula (1), we hope that the formula returns a larger value.
@@
Evidently, constraining the return value of the formula to achieve a minimum is equivalent to optimizing the value of $\mu_{\theta}$ in the formula. Conversely, constraining the return value to achieve a maximum is equivalent to minimizing the value of $\sigma_{\theta}$ in the formula. In other words, the probability-highlighting solver positions the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ for the reasoning problem through primary samples, while the standard deviation of the solution distribution is constrained by auxiliary samples.
@@
Based on the analysis in the previous section, auxiliary samples provided in reasoning tasks constrain the standard deviation of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$, while primary samples locate the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$. The Tine method generates more judicious auxiliary samples to further constrain the standard deviation. Although experimental results demonstrate the effectiveness of Tine, it remains uncertain whether Tine constrains the standard deviation too little or not enough. Furthermore, Tine based on adversarial learning also encounters all the issues associated with adversarial learning, such as unstable training processes and difficulties in determining the completeness of training. Given this series of challenges, this paper proposes the Funny (Framework Utilizing Neural Networks for Yielding Image Representation Distribution) method.
Our Funny method argues that instead of laboriously adopting adversarial learning strategies to make the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ provided by the solver closer to the correct solution distribution, it is more effective to directly specify the range of correct solutions or, alternatively, to explicitly define the details of the correct solution distribution. This paper will next analyze the setting and determination of correct solutions for image progressive pattern reasoning problems such as RPM and image clustering reasoning problems such as Bongard-Logo.
@@ VI-A Funny for RPM problems
This paper argues that by enforcing $bias_{i}$ to follow a standard normal distribution $N(0,1)$ and simultaneously enforcing all random sampling results, denoted as $\epsilon$, from a Gaussian distribution $N(center_{i},1)$ to assume the same reasoning identity, it can be ensured that $center_{i}$ stores at least the key attributes. This is because, when samples drawn from Gaussian distribution $N(center_{i},1)$ all share the same reasoning identity, $bias_{i}$, which follows a standard normal distribution $N(0,1)$, ceases to influence the reasoning, and $center_{i}$ takes full responsibility for the reasoning task, indicating that $center_{i}$ stores at least the key attributes at this point.
@@
Overall, Funny introduces three new MSE-based regression tasks for Valen. The green task aims to enforce the bias to conform to the standard normal distribution $N(0,1)$. The red task aims to maximize the mutual information between center+bias and $x_{i}$. The yellow task aims to enforce the output of the Funny decoder to conform to the distribution of $x_{i}$.
VI-B Funny, or not Funny
What is of paramount importance is that, potentially due to the specificity of MSE loss or unknown reasons arising from forcibly combining regression tasks with reasoning tasks, a decoder with a conventional structure cannot bear the aforementioned theoretical framework. If we apply a deconvolution structure to the backbone of the Funny decoder, its optimization becomes extremely challenging, making it difficult to achieve satisfactory regression results and supervise the mutual information between $center_{i}+bias_{i}$ and $x_{i}$. Alternatively, if we replace the Funny decoder with a ViT structure symmetric to the representation extractor, although we obtain excellent regression results, it leads to a decrease in the normal reasoning accuracy of Valen instead of an increase.
Unfortunately and unfunny, there is a disconnect between theory and practice. This paper speculates that this may be due to the impact of the regression task in Funny on the parameter optimization of the reasoning task in Valen. An attempt was made to linearly search for the scaling coefficient between the MSE-based regression task introduced by Funny and the CE-based reasoning task in Valen, but it did not yield an effective solution. Anyway, this paper does not intend to bridge this gap between theory and practice by adjusting the scaling coefficients among the various terms in the loss function.

### v2

# Funny-Valen-Tine: Planning Solution Distribution Enhances Machine Abstract Reasoning Ability

@@ Abstract
The importance of visual abstruct reasoning problems in the field of image processing is beyond words. Both Bongard-Logo problems and Raven’s Progressive Matrices (RPM) belong to the domain of visual abstract reasoning tasks, with Bongard-Logo categorized as image clustering reasoning and RPM involving image progression pattern reasoning. This paper introduces a novel baseline model, Valen, which falls under the umbrella of probabilistic highlighting models. Valen demonstrates remarkable performance in solving both RPM and Bongard-Logo problems, offering a versatile solution for these reasoning tasks. Our investigation extends beyond the application of Valen, delving into the underlying mechanisms of probability-highlighting solvers when tackling visual abstract reasoning problems. We revisit how these solvers handle instances in tasks like RPM and Bongard-Logo, realizing that they approximate the solution to each reasoning problem instance as a distribution delineated by primary and auxiliary samples. This led us to propose that the learning objective of probability-highlighting solvers is not the distribution of correct solutions but rather a distribution defined by both primary and auxiliary samples. To bridge the discrepancies, we introduced the Tine method, an adversarial learning-based approach designed to assist Valen in estimating a solution distribution closer to the correct one. However, Tine’s adversarial framework also introduces issues such as unstable training. Reflecting on Tine, we propose modeling the sample distribution of reasoning problems as a mixture of Gaussian distributions, leading to the development of the Funny method. Funny effectively models the sample distribution within reasoning problems as a mixture of Gaussian distributions, enabling Valen to more directly capture the true form of the correct solution distribution. Furthermore, we designed the SBR method to model the distribution of progressive patterns representation as a mixture of Gaussian distributions. Overall, this paper posits that the key to enhancing the ability of solvers to address visual abstract reasoning problems lies in explicitly planning the distribution of solutions to conform more closely to the distribution of correct solution.
@@
Abstract reasoning, Raven’s Progressive Matrices, Bongard-Logo, Adversarial Learning, Gaussian Mixture Model.
@@ I-A RAVEN and PGM
RPM[14], or Raven’s Progressive Matrices, is a classic visual reasoning intelligence test typically presented as a 3x3 matrix. It requires test-takers to observe and analyze visual features such as shape, size, direction, position, and their relationships within the problem statement, identify patterns or rules, and select an answer from given option pool that completes the matrix. Due to its non-verbal nature and value in assessing general intelligence, RPM is widely used in psychology, education, neuroscience, and other fields, serving as a tool for evaluating intelligence levels and studying human cognitive processes, intellectual development, and brain function. With the advancement of AI technology, researchers are exploring computer algorithms to solve RPM problems, albeit facing challenges. RAVEN[15] and PGM[16] are classic examples of RPM problems, with Figure 1 illustrating these two types of problems respectively.
@@ II-A RPM solvers and Bongard-Logo solvers
There are numerous excellent solvers for RPM and Bongard-Logo problems, among which CoPINet[19], LEN+teacher[18], DCNet[20], NCD[22], SCL[23], SAVIR-T[24], PrAE[25], NVSA[27], ALANS[26], RS-TRAN[28] and CRAB[29] are notable RPM solvers, while SBPDL[30], PMoC[31] and Triple-CFN[32] are outstanding Bongard-Logo solvers. Since the method proposed in this paper is relatively independent and does not have a strong dependency on these existing works, the details about them will not be elaborated further.
@@ III Methodology
In this paper, we designed a general solver, Valen, for both RPM and Bongard-Logo problems. This paper proposes that to enhance the ability of solvers to address visual abstract reasoning problems, the distribution of solutions provided by the solver should be explicitly planned to better conform to the distribution of correct solutions. Therefore, we developed three accompanying methods: Funny, Tine, and SBR. These three methods all aim to explicitly plan the distribution of solutions generated by Valen.
Here we clarify a definition: in this paper, “primary samples” refer to those that align with the core patterns in reasoning problems, like correct RPM options or primary group samples in Bongard-Logo instances. Conversely, “auxiliary samples” are those that deviate from these patterns, such as incorrect RPM options or auxiliary group samples in Bongard-Logo problems.
@@ IV Valen: a new baseline model for RPM and Bongard-Logo problems
In this section, we introduce Valen (Visual Abstraction Learning Network), our baseline model designed to address both RPM, an image progressive pattern reasoning problem, and Bongard-Logo, an image clustering reasoning problem.
@@ IV-A Valen targeting RPM problem
The RPM problem challenges participants to complete a 3×3 matrix of images by choosing answers from a pool of options. A correct answer is one that, when placed in the matrix, maintains a consistent and specific image progression pattern that can be deduced from the incomplete matrix, referred to as the problem statement[36].
@@ IV-A1 Representation extraction of matrix and enumeration of incomplete matrix
Subsequently, these multi-viewpoint representations $\{r_{i}|i\in[1,9]\}_{v}|_{v=1}^{V}$ will be enumerated by Valen to generate all possible incomplete representations matrix. The enumeration results are illustrated in the Figure 5.
@@
We denote these enumerated multi-viewpoint incomplete representation matrices as $\{r_{ni}|n\in[1,9],i\in[1,9],i\neq n\}_{v}|_{v=1}^{V}$, where $n$ represents the index of the incomplete representation matrix, and $i$ still represents the index of the representation within the incomplete matrix. Notably, the last item of the enumeration is the matrix with a incomplete option position, denoted as $\{r_{9i}|n\in[1,9],i\in[1,8]\}_{v}|_{v=1}^{V}$.
@@ IV-A2 Extraction of progressive patterns from incomplete matrix
The figure illustrates the process of the Transformer-Encoder handling $\{r_{9i}|i\in[1,8]\}_{v}$, which represents the last item of the enumeration results of incomplete representation matrices from a specific viewpoint $v$. The process of handling other incomplete representation matrices $\{r_{ni}|n\in[1,8],i\in[1,9],i\neq n\}_{v}$ is similar to the one shown in the figure, with no essential differences.
@@ V Tine, a Method for Advancing the Solution Distribution Boundary
Therefore, we can interpret these solvers as a conditional probability distributions obeyed by abstract reasoning problems, or understand them as providing distributions of solutions to abstract reasoning problems. If we use RPM solvers as examples, then these solvers can be abstracted as a conditional probability distribution, denoted as $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. (Accordingly, when Valen is applied to Bongard-Logo problems, it can be abstracted as a conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{6};\theta)$). In this context, the two conditions of the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ are the statement of the reasoning problem $\{x_{i}\}_{i=1}^{8}$ and the optimizable parameters $\theta$ of the solver. Here, the $x_{\alpha}$ refers to the option that is pending verification. This conditional distribution implies that, when given a problem statement $\{x_{i}\}_{i=1}^{8}$ and an option $x_{\alpha}$, it is possible to calculate the probability of $x_{\alpha}$ being a solution to $\{x_{i}\}_{i=1}^{8}$.
@@
The optimization of probability-highlighting solver can be regarded as optimizing the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. A wide range of loss functions can be used for optimization, but they all share the same objective: to assign high probabilities to the primary samples in the probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ and low probabilities to the auxiliary options. Such an optimization objective makes us realize that the optimization process of the probability-highlighting solver is not aimed at fitting the distribution of correct solutions to visual abstract reasoning problems, but rather is dedicated to fitting a distribution jointly delimited by primary and auxiliary samples.
@@
To simplify the explanation of our findings, we assume that the conditional probability distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ follows a Gaussian distribution. Consequently, we can explicitly identify the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ as the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$. Then, the negative logarithm of the probability density function of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$ can be expressed as follows:
$-\log(P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta}))=\frac{{(x_{a}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(|\,\sigma_{\theta}|\,)$ (1)
@@
Evidently, we have omitted the constant term $-1/2\log(2\pi)$. In this formula, $\mu_{\theta}$ and $\sigma_{\theta}$ respectively represent the mean and standard deviation of $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. Since we have defined the form of distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ as a Gaussian distribution, we consequently abstract the optimizable parameters $\theta$ into optimizable mean $\mu_{\theta}$ and standard deviation $\sigma_{\theta}$.
@@
Optimizing the distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\mu_{\theta},\sigma_{\theta})$ using the loss function is equivalent to optimizing the values of $\mu_{\theta}$ and $\sigma_{\theta}$. Our optimization objective is that when primary samples (the correct options in an RPM problem, or the samples within the main group in a Bongard-Logo problem) are plugged into formula (1), we expect to obtain a smaller value. Conversely, when auxiliary samples (the incorrect options in an RPM problem, or the samples within the auxiliary group in a Bongard-Logo problem) are plugged into formula (1), we hope that the formula returns a larger value.
@@
Evidently, constraining the return value of the formula to achieve a minimum is equivalent to optimizing the value of $\mu_{\theta}$ in the formula. Conversely, constraining the return value to achieve a maximum is equivalent to minimizing the value of $\sigma_{\theta}$ in the formula. In other words, the probability-highlighting solver positions the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ for the reasoning problem through primary samples, while the standard deviation of the solution distribution is constrained by auxiliary samples.
@@
Based on the analysis in the previous section, auxiliary samples provided in reasoning tasks constrain the standard deviation of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$, while primary samples locate the mean of the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$. The Tine method generates more judicious auxiliary samples to further constrain the standard deviation. Although experimental results demonstrate the effectiveness of Tine, it remains uncertain whether Tine constrains the standard deviation too little or not enough. Furthermore, Tine based on adversarial learning also encounters all the issues associated with adversarial learning, such as unstable training processes and difficulties in determining the completeness of training. Given this series of challenges, this paper proposes the Funny (Framework Utilizing Neural Networks for Yielding Image Representation Distribution) method.
Our Funny method argues that instead of laboriously adopting adversarial learning strategies to make the solution distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8};\theta)$ provided by the solver closer to the correct solution distribution, it is more effective to directly specify the range of correct solutions or, alternatively, to explicitly define the details of the correct solution distribution. This paper will next analyze the definition of correct solutions for image progressive pattern reasoning problems such as RPM and image clustering reasoning problems such as Bongard-Logo.
@@ VI-A Funny for RPM problems
This paper argues that by enforcing $bias_{i}$ to follow a standard normal distribution $N(0,1)$ and simultaneously enforcing all random sampling results from a Gaussian distribution $N(center_{i},1)$ to assume the same reasoning identity, it can be ensured that $center_{i}$ stores at least the key attributes. This is because, when samples drawn from Gaussian distribution $N(center_{i},1)$ all share the same reasoning identity, $bias_{i}$, which follows a standard normal distribution $N(0,1)$, ceases to influence the reasoning, and $center_{i}$ takes full responsibility for the reasoning task, indicating that $center_{i}$ stores at least the key attributes at this point.
@@
Our design intention for Funny is quite clear. We hope that, with the assistance of Valen, Funny can compel $center_{i}$ to store more key attributes related to reasoning, while $bias_{i}$ accumulates more non-key attributes unrelated to reasoning that could potentially interfere with it. By doing so, we aim to model the distribution of multi-viewpoint representations, generated by the $center_{i}+bias_{i}$, as a mixed Gaussian distribution that aids Valen in its reasoning process. The relationship between Funny and Valen is mutually beneficial and symbiotic. Although subsequent experiments demonstrated that Funny is a significantly effective method, given the fact that obtaining disentangled representations under unsupervised learning is not achievable [40], this paper holds a pessimistic attitude towards the extent of Funny’s decoupling of key and non-key attributes.
VI-B Funny not funny
What is of paramount importance is that, a Funny decoder with a conventional structure cannot bear the aforementioned theoretical framework. If we apply a deconvolution structure to the backbone of the Funny decoder, its optimization becomes extremely challenging, making it difficult to achieve satisfactory regression results and supervise the mutual information between $center_{i}+bias_{i}$ and $x_{i}$. Alternatively, if we replace the Funny decoder with a ViT structure symmetric to the representation extractor, although we obtain excellent regression results, it leads to a decrease in the normal reasoning accuracy of Valen instead of an increase.
Unfortunately and unfunny, there is a gep between theory and practice. This paper speculates that this may be due to the impact of the regression task in Funny on the parameter optimization of the reasoning task in Valen. If we choose the normal ViT as the Funny decoder, the introduced regression task will be completed before the Valen’s reasoning task, because the optimization speed of the regression task is faster than that of the reasoning task at this time. We posit that the premature completion of the regression task may result in the $center_{i}$ storing more non-key attribute. Because we have set up the regression task from $center_{i}+\epsilon$ to $x_{i}$. This paper speculates that this might be the issue: when the regression task approaches completion prematurely, it may lead to $center_{i}$ falling into the gradient vanishing zone, subsequently rendering it difficult for $center_{i}$ to make further optimizations for the reasoning task. The non-key attribute stored in $center_{i}$ can disrupt the reasoning process of Valen. An attempt was made to linearly search for the scaling coefficient between the MSE-based regression task introduced by Funny and the CE-based reasoning task in Valen, but it did not yield an effective solution. Anyway, this paper does not intend to bridge this gap between theory and practice by adjusting the scaling coefficients among the various terms in the loss function.